# Symmetry Breaking and Mie-tronic Supermodes in Nonlocal Metasurfaces

### v2

@@ Abstract
Symmetry breaking in metasurfaces is typically used as a means of weakening optical confinement, e.g. transforming infinite $Q$ bound states in the continuum into high $Q$ quasibound states. Within the Mie-tronics framework, we show that symmetry breaking can instead enhance light trapping by strengthening in-plane nonlocal coupling pathways. Through diffraction and multiple-scattering analyses, we demonstrate that diffractive bands and Mie-tronic supermodes originate from the same underlying Mie resonances but differ fundamentally in physical nature. Finite arrays exhibit $Q$-factor enhancement driven by redistributed radiation channels, reversing the trend predicted by infinite-lattice theory. We further show that controlled symmetry breaking opens new electromagnetic coupling channels, enabling polarization conversion in nonlocal metasurfaces. These findings establish a unified wave picture linking scattering and diffraction theories and outline design principles for multifunctional metasurfaces that exploit nonlocality for advanced light manipulation, computation, and emission control.
@@ I Introduction
Optical metasurfaces are rapidly evolving from fundamental science to practical technology, enabling compact platforms for imaging, sensing, energy harvesting, and information processing [Brongersma2025second]. Conventional metasurfaces comprise weakly interacting subwavelength elements, allowing a local description. Yet, recent analyses of their limits show that surpassing these constraints demands nonlocal design [Cai2020inverse, Shastri2023nonlocal, Yao2024nonlocal]. When inter-element coupling becomes strong, collective optical excitations emerge, giving rise to nonlocal metasurfaces governed by coherent, system-scale modes. Because these modes are typically excited by plane waves, they can often be interpreted through diffraction theory, earning the name diffractive nonlocal metasurfaces [Overvig2022diffractive]. Understanding such systems benefits from revisiting the classical diffraction theories that first distinguished local from nonlocal behavior.
@@
In his 1907 “Dynamical Theory of Gratings”, Rayleigh extended Fresnel’s scalar treatment of reflection and diffraction to periodic metallic surfaces. He explicitly noted that Fresnel’s local approximation fails when the grating period approaches the wavelength, observing that the recesses “act as resonators” [Rayleigh1907dynamical]. Rayleigh thus recognized, decades ahead of modern electromagnetic modeling, that resonance fundamentally alters diffraction and that a “more strictly dynamical theory” was required. Fano later provided this framework [Fano1941theory], introducing coupling among diffracted and evanescent orders to describe how these “Rayleigh cavities” form quasi-stationary surface waves that interfere with the continuum, giving rise to asymmetric Fano resonances. The Rayleigh–Fano formalism laid the foundation for modern numerical solvers such as the rigorous coupled-wave analysis (RCWA) or Fourier modal method [Liu2012s4], which we use here to calculate nonlocal diffraction bands in ideal, infinitely periodic metasurfaces.
@@
The design of nonlocal metasurfaces often begins with idealized infinite arrays, where symmetry breaking within the unit cell serves as a versatile tool for tailoring the dispersion and radiation properties of diffractive bands [Nguyen2018symmetry, Overvig2020selection, You2024resonance, Barreda2025bound]. Engineering such diffractive bands has recently fueled advances in analog optical computing and wave-based information processing [Guo2018photonic, Kwon2018nonlocal, Komar2021edge, Liu2024edge]. In practical implementations, however, metasurfaces are finite—a feature that has drawn growing attention [Ho2024finite, Karavaev2025emergence, Chen2025observation, Zhang2025vortex]. While compactness makes them appealing for integration, it simultaneously invalidates the assumption of infinite periodicity inherent to diffraction theory. As the array size decreases, translational symmetry is broken and the momentum conservation underlying the Fourier modal method no longer holds. The resulting collective resonances become discrete and strongly dependent on the global geometry of the structure. In this finite regime, multiple scattering and modal hybridization provide a more accurate physical description [Lagendijk1996resonant, Hoang2025collective].
@@
At the microscopic level, these collective effects originate from interactions among Mie modes—the localized resonances of individual dielectric or metallic elements—which constitute the building blocks of Mie-tronics [Rybin2024metaphotonics]. Beyond governing the response of isolated particles [Li2024machine, Barati2025toward], the coupling and interference of Mie modes across arrays yield bonding and antibonding supermodes. These supermodes can enhance device performance and enable entirely new metadevice concepts [Krasnok2012all, Babicheva2024mie, Ha2024optoelectronic, Mao2025lateral].
@@
In this work, we systematically investigate how symmetry breaking influences both diffractive and Mie-tronic supermodes. Together with the accompanying Letter [Hoang2025unconventional], we show that, contrary to expectations based on infinite-array diffraction bands, symmetry breaking in finite metasurfaces can enhance in-plane multiple scattering and thereby increase the quality factor ($Q$) of certain nonlocal modes. We further demonstrate that coupling among Mie modes produces additional bonding and antibonding supermodes beyond the Fourier-mode couplings of classical diffraction theory, expanding the functional landscape of nonlocal metasurfaces. Notably, while antibonding supermodes remain robust in the presence of a quartz substrate, bonding supermodes are suppressed. Finally, we show that distinct unit-cell geometries can support identical Mie-resonant supermodes, underscoring the robustness and design versatility of the Mie-tronic framework for both diffractive and finite-size nonlocal metasurfaces.
@@ II.1 Historical Origins of Mie-tronics
Over the following century, these contributions coalesced into the framework of generalized Lorenz–Mie theory, emphasizing the dual importance of intrinsic scatterer modes and the multipole content of excitation sources [Gouesbet2011generalized]. Advances in nanoscience have since extended scattering analysis to particles of arbitrary shape [Kildishev2025art], enabling entirely new classes of photonic devices [Dorodnyy2023mie]. This convergence of multipole physics and nanophotonics defines the emerging paradigm of Mie-tronics [Won2019into]—a framework that unifies classical scattering theory with modern metasurface design.
II.2 Diffraction and Scattering Frameworks
@@
II.3 Mie Coefficients as Building Blocks
@@
Spectral overlap among Mie modes thus lays the foundation for collective optical phenomena in arrays. When resonant unit cells are periodically coupled, their individual multipolar responses hybridize to form delocalized collective modes—Mie-tronic supermodes—in finite arrays. In the following section, we examine how these collective Mie-tronic supermodes connect to Bloch bands in infinite periodic systems.
II.4 Beyond Bloch Bands: Antibonding and Bonding Mie-tronic Supermodes
Figure 3: Mie-tronic origin of supermodes in finite arrays and their connection to Bloch bands in photonic crystals. (a) Purcell-factor spectrum for the configuration in the inset, revealing two well-separated bands (antibonding and bonding) with pronounced supermodes labeled $A_{1,2}$ and $B_{1}$. (b),(c) Magnetic ($H_{x}$) and electric ($E_{z}$) field distributions for the antibonding mode $A_{1}$; the inset shows the corresponding far-field radiation pattern. (d) Bloch bands associated with the antibonding and bonding supermodes in (a). (e),(f) Same as (b),(c), but for the bonding mode $B_{1}$; the inset highlights dominant in-plane radiation leakage.
Figure 3(a) presents the Purcell-factor spectrum for the configuration shown in the inset, where a magnetic dipole $m_{x}$ couples to a $5\times 5$ array of spheres. Two distinct bands appear, each containing well-defined supermodes labeled $A_{1}$, $A_{2}$, and $B_{1}$. The $A_{1}$ and $B_{1}$ modes correspond to the $\Gamma$- and $M$-point edges of the antibonding ($A$) and bonding ($B$) Bloch bands in Fig. 3(d), which emerge in the limit of an infinite array. To obtain the dispersions in Fig. 3(d), the dipole $m_{x}$ is positioned at the sphere center, and Bloch boundary conditions are applied along the $y$ and $z$ directions.
@@ II.4 Beyond Bloch Bands: Antibonding and Bonding Mie-tronic Supermodes
The antibonding and bonding characters of these modes are evident from their near- and far-field profiles. Figures 3(b,c) show the $A_{1}$ (antibonding) supermode, while Figs. 3(e,f) show the $B_{1}$ (bonding) mode. The in-plane electric-field ($E_{z}$) distributions reveal that $A_{1}$ exhibits nodes between adjacent spheres, whereas $B_{1}$ displays pronounced hotspots within the interparticle gaps. These hotspots signify stronger in-plane coupling and larger Purcell enhancement for the bonding mode. Such enhanced local fields enable efficient interaction with emitters or detectors oriented along the $z$ direction, suggesting on-chip architectures where Mie-tronic modes act as optical communication channels linking integrated transmitters and receivers—forming scalable photonic networks capable of emulating quantum many-body interactions. The topological properties of Mie modes, in this context, could underpin future classical and nonclassical simulators and light sources [Nielsen2010quantum, Berloff2017realizing, Sergeeva2025coherent, Deng2025enhancement].
A further distinction arises from the orientation of the magnetic moments. In $A_{1}$, the moments are aligned, leading to repulsive interactions; in $B_{1}$, they are anti-aligned, resulting in attractive coupling. Put differently, bonding supermodes extend their fields into the environment, whereas antibonding modes confine them within the unit cells. This difference explains the greater robustness of antibonding modes to perturbations introduced by a quartz substrate, as discussed in the next section.
The far-field radiation patterns [insets of Figs. 3(b,e)] provide complementary evidence. The bonding mode $B_{1}$ radiates mainly in-plane, while the antibonding mode $A_{1}$ emits strongly out of plane. Similar to the hole arrays discussed in the accompanying Letter [Hoang2025unconventional], the $A_{1}$ mode produces an out-of-plane vortex beam, confirming its collective MD origin. From an applications perspective, the in-plane confinement of $B_{1}$ makes it especially appealing for on-chip photonic devices that demand minimal out-of-plane losses.
@@
Nevertheless, the collective nature of Mie-tronic supermodes allows normal-incidence excitation in sufficiently large arrays [Hoang2025collective]. In Fig. 3(a), only three discrete supermodes appear for the $5\times 5$ array, but their number increases rapidly with array size. Retardation introduces additional supermodes whose $E_{z}$ distributions become symmetric about the $z$ axis, enabling coupling to normally incident plane waves. The discrete peaks in Fig. 3(a) arise from the radiative boundary conditions of finite arrays, capturing high-$Q$ supermodes inaccessible to conventional Bloch-wave analysis, which assumes in-plane periodicity and thus yields only continuous bonding and antibonding bands [Fig. 3(d)].
@@
These finite-array effects highlight how the interplay between collective coupling and structural symmetry profoundly reshapes resonance behavior—a theme we now pursue by examining how symmetry breaking governs the formation and quality factors of Mie-tronic supermodes.
III Impact of Symmetry Breaking on Mie-tronic Supermodes
@@
Symmetry breaking in unit cells has been extensively explored for engineering diffractive bands, yet its influence on collective resonances in finite metasurfaces remains far less understood. Here, we show that symmetry breaking not only modifies the coupling between Mie-tronic supermodes and external radiation but can also produce counterintuitive $Q$-factor enhancement—opposite to trends predicted by infinite-lattice Bloch theory. Moreover, the distinct topological characteristics of Mie modes provide a natural foundation for designing nonlocal metasurfaces with tailored confinement, coupling, and polarization properties.
III.1 Robustness of Antibonding Supermodes Against Symmetry Breaking
@@
Figure 4: Symmetry breaking enhances in-plane multiple scattering, preserves antibonding supermodes, and suppresses bonding counterparts in square and T-shaped arrays. (a) Purcell-factor spectrum for a magnetic dipole $m_{x}$ at the center of a $9\times 9$ array of square unit cells. The presence of a quartz substrate ($n\simeq 1.45$) suppresses the bonding mode $B_{1}$, while the antibonding mode $A_{1}$ remains nearly unaffected. (b),(c) Magnetic near-field distribution and vortex-like far-field pattern of $A_{1}$. (d)–(f) Same as (a)–(c) but for the T-shaped array. The far-field pattern highlights enhanced in-plane multiple scattering along in-plane channels (ICs) versus suppressed out-of-plane channels (OCs).
@@
Figure 4 summarizes the effects of in-plane and out-of-plane symmetry breaking on Mie-tronic supermodes. In-plane symmetry breaking is introduced by replacing square unit cells with T-shaped ones (same lattice vectors but reduced unit-cell symmetry), while out-of-plane breaking is implemented by adding a quartz substrate. For the free-standing square lattice, two well-separated resonant bands appear [Fig. 4(a)], corresponding to the antibonding ($A_{1}$) and bonding ($B_{1}$) supermodes at the band edge. The magnetic-field distribution of $A_{1}$ [Fig. 4(b)] confirms its MD nature, consistent with the sphere array in Fig. 3. Introducing a substrate strongly perturbs the bonding mode $B_{1}$, whose attractive field distribution extends into the environment and overlaps with substrate-supported channels. In contrast, the antibonding mode remains largely unaffected owing to its repulsive character, which minimizes near-field interaction with the substrate.
@@
III.2 $Q$ Enhancement via Symmetry Breaking
Figure 5: Symmetry breaking lowers the $Q$ of Bloch modes in infinite lattices but enhances the $Q$ of supermodes in finite metasurface arrays. (a) Antibonding Bloch bands above the light line for the four unit-cell types. (b) $Q$ of Bloch modes at $\Gamma$: symmetry breaking reduces the divergent $Q$ of the square unit to a finite value in the T-shaped unit, while sphere and hole lattices show much higher $Q$. (c) $Q$ of supermodes in finite arrays: symmetry breaking boosts the $Q$ of T-shaped arrays beyond that of square arrays for sizes $5\times 5$ to $33\times 33$, with both outperforming sphere and hole arrays.
Figure 5(a) shows the antibonding Bloch bands above the light line for the four unit-cell geometries. The calculated band-edge wavelengths agree well with the corresponding $A_{1}$ supermodes (the hole-array case is detailed in the accompanying Letter [Hoang2025unconventional]). Modifying the unit-cell geometry shifts the band edges and alters the dispersion; notably, the square and T-shaped lattices exhibit narrow spectral widths ($\sim$20 nm), nearly an order of magnitude smaller than that of the hole lattice (179 nm). This result highlights the capability of geometrical design to engineer band dispersion with high precision.

### v3

# Mie-tronics supermodes and symmetry breaking in nonlocal metasurfaces

@@ Abstract
It is usually believed that symmetry breaking in photonic systems leads to weaker optical confinement, such as in the case of metasurfaces when bound states in the continuum are replaced by quasi-bound states with lower quality factors ($Q$ factors). Here we show that symmetry breaking can instead enhance light trapping by strengthening in-plane nonlocal coupling pathways. We consider finite-size arrays of optical resonators supporting Mie resonances (a Mie-tronics platform) and employ diffraction and multiple-scattering analyses. We demonstrate that diffractive bands and Mie-tronics supermodes originate from the same underlying Mie resonances but differ fundamentally in their physical nature. Finite arrays exhibit $Q$-factor enhancement driven by redistributed radiation channels, and reversing the trends predicted by infinite-lattice theories. We reveal that controlled symmetry breaking opens new electromagnetic coupling channels, enabling polarization conversion in nonlocal metasurfaces. These novel findings establish a unified wave-physics platform linking both scattering and diffraction theories. Also, they outline the design principles for multifunctional metasurfaces that exploit nonlocality for advanced light manipulation, computation, and emission control.
@@ I Introduction
The field of optical metasurfaces is rapidly evolving from fundamental science to practical technologies, enabling compact platforms for imaging, sensing, energy harvesting, and information processing [Brongersma2025second]. Conventional metasurfaces comprise weakly interacting subwavelength elements, allowing a local description. Yet, recent analyses of their limits show that surpassing these constraints demands nonlocal design [Cai2020inverse, Shastri2023nonlocal, Yao2024nonlocal]. When inter-element coupling becomes strong, collective optical excitations emerge, giving rise to nonlocal metasurfaces governed by coherent, system-scale modes. Because these modes are typically excited by plane waves, they can often be interpreted through diffraction theory, earning the name diffractive nonlocal metasurfaces [Overvig2022diffractive]. Understanding such systems benefits from revisiting the classical diffraction theories that first distinguished local from nonlocal behavior.
@@
In his 1907 “Dynamical Theory of Gratings”, Rayleigh extended Fresnel’s scalar treatment of reflection and diffraction to periodic metallic surfaces. He explicitly noted that Fresnel’s local approximation fails when the grating period approaches the wavelength, observing that the recesses “act as resonators” [Rayleigh1907dynamical]. Rayleigh thus recognized, decades ahead of modern electromagnetic modeling, that resonance fundamentally alters diffraction and that a “more strictly dynamical theory” is required. Ugo Fano later provided this framework [Fano1941theory], introducing coupling among diffracted and evanescent orders to describe how these “Rayleigh cavities” form quasi-stationary surface waves that interfere with the continuum, giving rise to asymmetric Fano lineshapes. The Rayleigh–Fano formalism laid the foundation for modern numerical solvers such as the rigorous coupled-wave analysis (RCWA) or Fourier modal method [Liu2012s4], which are employed here to calculate nonlocal diffraction bands in ideal, infinitely periodic metasurfaces.
@@
The design of nonlocal metasurfaces often begins with idealized infinite arrays of resonators, where symmetry breaking within the unit cell serves as a versatile tool for tailoring the dispersion and radiation properties of diffractive bands [Nguyen2018symmetry, Overvig2020selection, You2024resonance, Barreda2025bound]. Engineering such diffractive bands has recently fueled advances in analog optical computing and wave-based information processing [Guo2018photonic, Kwon2018nonlocal, Komar2021edge, Liu2024edge]. In practical implementations, however, metasurfaces are finite—a feature that has drawn growing attention [Ho2024finite, Karavaev2025emergence, Chen2025observation, Zhang2025vortex]. While compactness makes them appealing for integration, it simultaneously invalidates the assumption of infinite periodicity inherent to diffraction theory. As the array size decreases, translational symmetry is broken and the momentum conservation underlying the Fourier modal method no longer holds. The resulting collective resonances become discrete and strongly dependent on a global geometry of the structure. In this finite regime, multiple scattering and modal hybridization provide a more accurate physical description [Lagendijk1996resonant, Hoang2025collective].
@@
At the microscopic level, the collective effects originate from interactions among Mie modes—the localized resonances of individual dielectric or metallic elements—which constitute the building blocks of Mie-tronics [Rybin2024metaphotonics]. Beyond governing the response of isolated particles [Li2024machine, Barati2025toward], the coupling of Mie modes across arrays yield bonding and anti-bonding supermodes. These supermodes can enhance device performance and enable entirely new metadevice concepts [Krasnok2012all, Babicheva2024mie, Ha2024optoelectronic, Mao2025lateral].
@@
In this work, we study systematically how symmetry breaking can affect both diffractive and Mie-tronics supermodes. Together with the accompanying Letter [Hoang2025unconventional], we show that, contrary to expectations based on infinite-array diffraction bands, symmetry breaking in finite metasurfaces can enhance in-plane multiple scattering and thereby increase the quality factor ($Q$) of certain nonlocal modes. We further demonstrate that coupling among Mie modes produces additional bonding and anti-bonding supermodes beyond the Fourier-mode couplings of classical diffraction theory, expanding the functional landscape of nonlocal metasurfaces. Notably, while anti-bonding supermodes remain robust in the presence of a quartz substrate, bonding supermodes are suppressed. Finally, we show that distinct unit-cell geometries can support identical Mie-resonant supermodes, underscoring the robustness and design versatility of the Mie-tronics framework for both diffractive and finite-size nonlocal metasurfaces.
@@ II.1 Historical origins of Mie-tronics
Over the following century, these contributions coalesced into the framework of generalized Lorenz–Mie theory, emphasizing the dual importance of intrinsic scatterer modes and the multipole content of excitation sources [Gouesbet2011generalized]. Advances in nanoscience have since extended scattering analysis to particles of arbitrary shape [Kildishev2025art], enabling entirely new classes of photonic devices [Dorodnyy2023mie]. This convergence of multipole physics and nanophotonics defines the emerging paradigm of Mie-tronics [Won2019into, kivshar_Mie]—a framework that unifies classical scattering theory with modern metasurface design.
II.2 Diffraction and scattering
@@
II.3 Mie modes as the building blocks
@@
Spectral overlap among Mie modes thus lays the foundation for collective optical phenomena in arrays. When resonant unit cells are periodically coupled, their individual multipolar responses hybridize to form delocalized collective modes—Mie-tronics supermodes—in finite arrays. In the following section, we examine how these collective Mie-tronics supermodes connect to Bloch bands in infinite periodic systems.
II.4 Beyond Bloch bands: Anti-bonding and bonding supermodes
Figure 3: Mie-tronics origin of supermodes in finite arrays and their connection to Bloch bands in photonic crystals. (a) Purcell-factor spectrum for the configuration in the inset, revealing two well-separated bands (anti-bonding and bonding) with pronounced supermodes labeled $A_{1,2}$ and $B_{1}$. (b),(c) Magnetic ($H_{x}$) and electric ($E_{z}$) field distributions for the anti-bonding mode $A_{1}$; the inset shows the corresponding far-field radiation pattern. (d) Bloch bands associated with the anti-bonding and bonding supermodes in (a). (e),(f) Same as (b),(c), but for the bonding mode $B_{1}$; the inset highlights dominant in-plane radiation leakage.
Figure 3(a) presents the Purcell-factor spectrum for the configuration shown in the inset, where a magnetic dipole $m_{x}$ couples to a $5\times 5$ array of spheres. Two distinct bands appear, each containing well-defined supermodes labeled $A_{1}$, $A_{2}$, and $B_{1}$. The $A_{1}$ and $B_{1}$ modes correspond to the $\Gamma$- and $M$-point edges of the anti-bonding ($A$) and bonding ($B$) Bloch bands in Fig. 3(d), which emerge in the limit of an infinite array. To obtain the dispersions in Fig. 3(d), the dipole $m_{x}$ is positioned at the sphere center, and Bloch boundary conditions are applied along the $y$ and $z$ directions.
@@ II.4 Beyond Bloch bands: Anti-bonding and bonding supermodes
The anti-bonding and bonding characters of these modes are evident from their near- and far-field profiles. Figures 3(b,c) show the $A_{1}$ (anti-bonding) supermodes, while Figs. 3(e,f) show the $B_{1}$ (bonding) mode. The in-plane electric-field ($E_{z}$) distributions reveal that $A_{1}$ exhibits nodes between adjacent spheres, whereas $B_{1}$ displays pronounced hotspots within the interparticle gaps. These hotspots signify stronger in-plane coupling and larger Purcell enhancement for the bonding mode. Such enhanced local fields enable efficient interaction with emitters or detectors oriented along the $z$ direction, suggesting on-chip architectures where Mie-tronics modes act as optical communication channels linking integrated transmitters and receivers—forming scalable photonic networks capable of emulating quantum many-body interactions. The topological properties of Mie modes, in this context, could underpin future classical and nonclassical simulators and light sources [Nielsen2010quantum, Berloff2017realizing, Sergeeva2025coherent, Deng2025enhancement].
A further distinction arises from the orientation of the magnetic moments. In $A_{1}$, the moments are aligned, leading to repulsive interactions; in $B_{1}$, they are anti-aligned, resulting in attractive coupling. Put differently, bonding supermodes extend their fields into the environment, whereas anti-bonding modes confine them within the unit cells. This difference explains the greater robustness of anti-bonding modes to perturbations introduced by a quartz substrate, as discussed in the next section.
The far-field radiation patterns [insets of Figs. 3(b,e)] provide complementary evidence. The bonding mode $B_{1}$ radiates mainly in-plane, while the anti-bonding mode $A_{1}$ emits strongly out of plane. Similar to the hole arrays discussed in the accompanying Letter [Hoang2025unconventional], the $A_{1}$ mode produces an out-of-plane vortex beam, confirming its collective MD origin. From an applications perspective, the in-plane confinement of $B_{1}$ makes it especially appealing for on-chip photonic devices that demand minimal out-of-plane losses.
@@
Nevertheless, the collective nature of Mie-tronics supermodes allows normal-incidence excitation in sufficiently large arrays [Hoang2025collective]. In Fig. 3(a), only three discrete supermodes appear for the $5\times 5$ array, but their number increases rapidly with array size. Retardation introduces additional supermodes whose $E_{z}$ distributions become symmetric about the $z$ axis, enabling coupling to normally incident plane waves. The discrete peaks in Fig. 3(a) arise from the radiative boundary conditions of finite arrays, capturing high-$Q$ supermodes inaccessible to conventional Bloch-wave analysis, which assumes in-plane periodicity and thus yields only continuous bonding and anti-bonding bands [Fig. 3(d)].
@@
These finite-array effects highlight how the interplay between collective coupling and structural symmetry profoundly reshapes resonance behavior—a theme we now pursue by examining how symmetry breaking governs the formation and quality factors of Mie-tronics supermodes.
III Impact of symmetry breaking on Mie-tronics supermodes
@@
Symmetry breaking in unit cells has been extensively explored for engineering diffractive bands, yet its influence on collective resonances in finite metasurfaces remains far less understood. Here, we show that symmetry breaking not only modifies the coupling between Mie-tronics supermodes and external radiation but can also produce counterintuitive $Q$-factor enhancement—opposite to trends predicted by infinite-lattice Bloch theory. Moreover, the distinct topological characteristics of Mie modes provide a natural foundation for designing nonlocal metasurfaces with tailored confinement, coupling, and polarization properties.
III.1 Robustness of anti-bonding supermodes against symmetry breaking
@@
Figure 4: Symmetry breaking enhances in-plane multiple scattering, preserves anti-bonding supermodes, and suppresses bonding counterparts in square and T-shaped arrays. (a) Purcell-factor spectrum for a magnetic dipole $m_{x}$ at the center of a $9\times 9$ array of square unit cells. The presence of a quartz substrate ($n\simeq 1.45$) suppresses the bonding mode $B_{1}$, while the anti-bonding mode $A_{1}$ remains nearly unaffected. (b),(c) Magnetic near-field distribution and vortex-like far-field pattern of $A_{1}$. (d)–(f) Same as (a)–(c) but for the T-shaped array. The far-field pattern highlights enhanced in-plane multiple scattering along in-plane channels (ICs) versus suppressed out-of-plane channels (OCs).
@@
Figure 4 summarizes the effects of in-plane and out-of-plane symmetry breaking on Mie-tronics supermodes. In-plane symmetry breaking is introduced by replacing square unit cells with T-shaped ones (same lattice vectors but reduced unit-cell symmetry), while out-of-plane breaking is implemented by adding a quartz substrate. For the free-standing square lattice, two well-separated resonant bands appear [Fig. 4(a)], corresponding to the anti-bonding ($A_{1}$) and bonding ($B_{1}$) supermodes at the band edge. The magnetic-field distribution of $A_{1}$ [Fig. 4(b)] confirms its MD nature, consistent with the sphere array in Fig. 3. Introducing a substrate strongly perturbs the bonding mode $B_{1}$, whose attractive field distribution extends into the environment and overlaps with substrate-supported channels. In contrast, the anti-bonding mode remains largely unaffected owing to its repulsive character, which minimizes near-field interaction with the substrate.
@@
III.2 $Q$-enhancement via symmetry breaking
Figure 5: Symmetry breaking lowers the $Q$ of Bloch modes in infinite lattices but enhances the $Q$ of supermodes in finite metasurface arrays. (a) anti-bonding Bloch bands above the light line for the four unit-cell types. (b) $Q$ of Bloch modes at $\Gamma$: symmetry breaking reduces the divergent $Q$ of the square unit to a finite value in the T-shaped unit, while sphere and hole lattices show much higher $Q$. (c) $Q$ of supermodes in finite arrays: symmetry breaking boosts the $Q$ of T-shaped arrays beyond that of square arrays for sizes $5\times 5$ to $33\times 33$, with both outperforming sphere and hole arrays.
Figure 5(a) shows the anti-bonding Bloch bands above the light line for the four unit-cell geometries. The calculated band-edge wavelengths agree well with the corresponding $A_{1}$ supermodes (the hole-array case is detailed in the accompanying Letter [Hoang2025unconventional]). Modifying the unit-cell geometry shifts the band edges and alters the dispersion; notably, the square and T-shaped lattices exhibit narrow spectral widths ($\sim$20 nm), nearly an order of magnitude smaller than that of the hole lattice (179 nm). This result highlights the capability of geometrical design to engineer band dispersion with high precision.